OpenStudy (anonymous):

Find 22E i=9 (10i+3) Sorry, I'm not quite sure how to post this equation. I tired solving it but I got the wrong answer (14,532). Could anyone help me out? Thank you(:

OpenStudy (phi):

can you use the equation editor (see the button on the lower left of the input area) also, does i mean sqrt(-1) ? what is E ?

OpenStudy (anonymous):

\[\sum_{9}^{22} 10i+3\]

OpenStudy (mertsj):

\[a _{n}=a _{1}(n-1)d\]

OpenStudy (mertsj):

\[a _{14}=a _{9}(14-1)(d)=93(13)(10)=12090\]

OpenStudy (anonymous):

I see how you got your answer but that's not one of my choices. 2,054 2,119 2,142 2,212 Also, my lesson said that I should use this equation Sn=n/2(a1+ai) but when I used that, I got the wrong answer

OpenStudy (anonymous):

I'm sorry, replace the n's with i's. So, Si=i/2(a1+ai)

OpenStudy (mertsj):

\[a _{14}=a _{9}+(14−1)(d)=93+(13)(10)=223\]

OpenStudy (mertsj):

I missed the plus sign before. Now use the formula for the sum. The ninth term is 93 and the 22th term is 223 so the number of terms is 14

OpenStudy (mertsj):

Yes. There is another fine way.

OpenStudy (mertsj):

Why don't you post it...then the asker can have his/her choice

OpenStudy (mertsj):

I guess you mean: \[a _{n}=a _{0}+(n-1)(10)\]

OpenStudy (mertsj):

Oh. So first term is 0?

OpenStudy (loser66):

this is mine.

OpenStudy (anonymous):

Oh I see, my lesson didn't explain that very well lol. Thank you, both of you! (:

OpenStudy (loser66):

mine makes sense to you? :)

OpenStudy (anonymous):

It took me a little bit to understand the order of your work and how you did some things but essentially I understand.

OpenStudy (loser66):

my prof asked me to do this to make sure that I understand how to get the general form of sequence whose the next number doesn't relate to the previous one. but somehow unclear . Ok, anyway, at the end up, you can get it and i don't run away. hehehe

OpenStudy (anonymous):

Thanks for your help!

OpenStudy (loser66):

ok, just give out opinion,